23- ter 01 Problems Help Sa Milden Company is a merchandiser that plans to sell 31,000 units during the next quarter at a selling price of $55 per unit. The company also gathered the following cost estimates for the next quarter Cost of good sold Advertising expense Sales commissions Shipping expense Administrative salaries $175,000 per quarter $60,000 per quarter $5.00 per unit sold $85,000 per quarter $9,500 per quarter $55,000 per quarter Depreciation expense Required 1. Prepare a contribution format income statement for the next quarter 2. Prepare a traditional format income statement for the next quarter. Complete this question by entering your answers in the tabs below Required 1 Required 2 Prepare a contribution format income statement for the next quarter. Milden For the Next Quarter
Chee n format i
Required 1 Required 2 Prepare a contribution format income statement for the next quarter. Milden Company Income For the Next Quarter es Total fixed expenses ot 4 Next
06- Chapter 01 Problems 4 For the Next Qu Total selling and a Required 1
Answer #1

CONTRIBUTION FORMAT INCOME STATEMENT

Sales (31000 x $55) 1705000
Less: Variable Costs -257300
Contribution Margin 1447700
Less: Fixed Costs -384500
EBIT 1063200

Variable Costs: Sales Commission + Variable Shipping Expenses = (1705000*6/100)+($5*31000) = $257300

Fixed Costs = 175000+60000+85000+9500+5500 = $384500

TRADITIONAL FORMAT INCOME STATEMENT

Sales 1705000
Less: COGS (25*31000) -775000
Gross Profit 930000
Less: Selling & Administrative Expenses
Variable(1075000*6/100) + (5*31000) -219500
Fixed (+175000+60000+85000+9500) -105000
Less: General Fixed Costs (Depreciation) -55000
Operating Income 550500